I am trying to teach my students quadriatic equations but I need to understand them myself. We use the Plato program but I keep coming up with the wrong answer. Can you give me the steps.... in laymans terms. I am NOT a math brained person. Two of the problems that have given us trouble are:
1.) 24 + 11y + y2 = 0
and
2.) 5y - 14 = -y2

Thanks

Answer by stanbon(75887)

I am trying to teach my students quadriatic equations but I need to understand them myself. Two of the problems that have given us trouble are:
Quadratic equations are solved by: 1) graphing ; 2) factoring ; and
3) quadratic formula.
---------------------------
1.) 24 + 11y + y^2 = 0
The formula is x = [-b +- sqrt(b^2 - 4ac)]/(2a)
In your problem, a = 1, b = 11, c = 24
So x = [-11 +- sqrt(11^2 - 4*1*24)]/2
x = [-11 +- sqrt(121 - 96)]/2
x = [-11 +- sqrt(25)]/2
x = [-11 +- 5]/2
x = -8 or x = -3
--------------------
and
2.) 5y - 14 = -y^2
Rearrange to get:
y^2 + 5y - 14 = 0
Factor:
(y+7)(y-2) = 0
Since the product is zero, one of the factors must be zero:
y = -7 or y = 2
==========================
Cheers,
Stan H.